Fermi National Laboratory
 

Operation Summary for September 14, 2011 - September 16, 2011

 

Wednesday September 14

The day shift began with the Tevatron (TEV) cryo systems recovering, with Operations monitoring the Recycler stash (205.2E10), with the Antiproton Source (Pbar) stacking (34.8E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next TEV store.

During the day shift, Cryo System personnel reported at 8:41 AM that the TEV sector E2 wet engine belts were slipping; they had them tightened and the engine back online within 30 minutes. The Booster West Anode Power Supply tripped off at 8:56 AM; a RF person investigated and had the supply back on within 15 minutes. Operators had current on the TEV bus and started ramping the TEV, but it tripped off at 11:21 AM. Operators investigated and reported a mechanical dump switch failure at E3; an EE Support person replaced its DC breaker. Operations put the TEV through a wet squeeze at 1:32 PM. The TEV suffered a low beta quench at 2:19 PM due to bad beam positions. The quench also caused a TEV sector C2 dump failure; EE Support personnel began repairs. Operators rebooted MECAR at 2:50 PM.

During the evening shift, Operators put the TEV through a dry squeeze at 5:43 PM. Operations established a 450 GeV 3x3 study store at 8:19 PM. The Booster low level RF room air conditioning unit failed; Booster and FESS personnel investigated and turned on the standby unit.

Thursday September 15

The midnight shift began with Operations monitoring a 3x3 study store and the stash (196.2E10), with Pbar stacking (234.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, at 2 AM, the Booster beam permit tripped off due to the loss of communication with a Booster RF station (BRF10). At 3:07 AM, operators received vacuum alarms from the TEV and Switchyard; operators rebooted the TEVAC system, but that didn’t help. Operators, searching for the Booster communication problem, found a rack monitor (IRM ) module off; operators called in Controls personnel who, by 3:30 AM, had the rack monitor card repaired. The I- Source efficiency had dropped off after the Booster trip. Operators consulted with Linac personnel and were able to get the efficiency back up. A Linac RF station (LRF1) tripped off several times; it reset. Operations disabled beam to NuMI at 3:30 AM in preparation for access in the morning to change out the target. Operations terminated the TEV 3x3 study store at 4:59 AM. Operators put the TEV through a dry squeeze. Operations established another TEV 3x3 study store at 6:20 AM. At 7:23 AM, the TEV 3x3 store aborted due to a low beta spool quadrupole magnet power supply (BQ0) trip on an over-voltage fault; EE Support personnel investigated. Operators prepared D0 for an access at 7:43 AM.

During the day shift, Operations received the Operational Readiness Clearance for the next Meson FTBF experiment T-922 at 8:37 AM. EE Support personnel reported at 9:57 AM that they had repaired C:BQ0. D0 came out of access at 10:01 AM. Operators put the TEV through a dry squeeze at 12:03 PM and a wet squeeze at 12:43 PM. At 1:22 PM, the TEV suffered a quench during a wet squeeze. Operations established beam to T-922 at 1:33 PM.

During the evening shift, operators put the TEV through a dry squeeze at 4:06 PM and a wet squeeze at 4:45 PM. Operations disabled beam to T-922 around 5:30 PM. Operations established at 3x3 study store at 6:18 PM. Operators reported short downtimes due to problems with LRF3 and LRF5.

Friday September 16

The midnight shift began with Operations monitoring a 3x3 study store and the stash (202E10), with Pbar stacking (260.8E10), with NuMI off for a target change, and with MiniBooNE taking beam.

During the midnight shift, Operations terminated the 3x3 study store at 1:40 AM and put the TEV through a dry squeeze. FIRUS reported network communication problems at 2:25 AM; operators called in a Controls person to investigate. Operations established a 3x3 study store at 3:02 AM. Operators reported at 4:37 AM that the Pbar Debuncher kicker (D:EKIK) had stopped putting out voltage; operators investigated and then called in a Kicker Power Supply person. At 6:48 AM, a Linac quadrupole bulk power supply tripped off; operators called in Linac technicians. The Kicker Power Supply person investigating D:EKIK said that the 80kV power supply was bad and needed replacing; he began the work.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow Booster personnel to conduct a study. The TEV has completed its run of low energy study stores.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/16/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory